6. The two blocks, m = 10kg and m = 50 kg are free to move as shown. The coefficient of static friction between the blocks Is 0.5 and there is no friction between M and the ground. A minimum horizontal force F is applied to hold m against M that is equal to – ( plz explain the solution in detail)

6.     The two blocks, m = 10kg and m = 50 kg are free to move as shown. The coefficient of static friction between the blocks Is 0.5 and there is no friction between M and the ground. A minimum horizontal force F is applied to hold m against M that is equal to – ( plz explain the solution in detail)


 


7jul10-img1.jpg

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

Let F be exerted on m. let there be no slipping,acceleration of two blocks at rest with respect to each other=F/(M+m)

force exerted by m on M=MF/(M+m)=normal force.,downward force of gravity on m=mg,maximum friction=(1/2)(MF/(M+m))

thus MF/(2(M+m))≥mg, F≥2mg(1+(m/M))